# Can the diffeomorphism and Gauss constraints be holonomy corrected in the deformed algebra approach to modified gravity?

### v1

@@ Abstract
Deforming the algebra of constraint is a well known approach to effective loop quantum cosmology. More generally, it is a consistent way to modify gravity from the Hamiltonian perspective. In this framework, the Hamiltonian (scalar) constraint is usually the only one to be holonomy corrected. As a heuristic hypothesis, we consider the possibility to also correct the diffeomorphism and Gauss constraints. It is shown that it is impossible to correct the diffeomorphism constraint without correcting the Gauss one, while maintaining a first-class algebra. However, if all constraints are corrected, the algebra can be closed. The resulting differential equations to be fulfilled by the corrections (of the background and of the perturbations) are derived.
@@ I Introduction
In the loop view, quantum geometry corrections are effectively taken into account by replacing the Sen-Ashtekar-Barbero connection by a generalized function which encodes the holonomy along a graph, the so-called holonomy correction. Following recent works – in particular remarks made in [19, 20, 21, 22] – we implement this correction into the classical algebra of general relativity, without specifying its explicit form a priori. Then, we ensure the consistency of the model by requiring the closure of the algebra of constraints – that is in making sure that each Poisson bracket between constraints is proportional to a constraint, or a combination of constraints: the evolution of a physical solution yields another physical solution. This approach actually extends far beyond its initial purpose and offers a powerful framework for studying modified gravity.
@@
At the end of the day, one should obtain a first class Poisson algebra and the subtle cancellation game between anomalies and counterterms could lead to restrictions on the allowed forms of the correction functions. This might have observable consequences as the Mukhanov-Sasaki equation for gauge-invariant perturbations is expected to be modified accordingly. Not to mention a possible disappearance of time [23, 24] around the quantum bounce [25].
Usually, the Hamiltonian constraint is the only one to receive corrections. Although there are reasons for this [26], we investigate here, at the heuristic level, the possibility that the diffeomorphism and Gauss constraints may themselves receive corrections. We begin by focusing on the correction of the diffeomorphism constraint alone. We then proceed by considering the full set of corrections affecting all the constraints that define the theory. The differential equations to be fulfilled by the corrections are investigated in details. We finally conclude with some remarks on possible future studies.
II Correcting the Hamiltonian and diffeomorphism constraints only
@@ II.1 Expression of the constraints
In this work, in addition to the Hamiltonian constraint, the diffeomorphism constraint is also corrected by a function $h(\mathfrak{c},\mathfrak{p})$. Following the usual procedure, three new counter-terms $\xi_{i}(\mathfrak{c},\mathfrak{p})$, $i=1,2,3$, are then introduced. The geometrical constraint reads :
@@
with $N^{a}={\bf{N}}^{a}+\delta N^{a}$ the shift function, $E_{i}^{a}=\mathfrak{p}\delta_{i}^{a}+\delta E^{a}_{i}$ the densitized triads and $K_{a}^{i}=\mathfrak{c}\delta^{i}_{a}+\delta K^{i}_{a}$ their canonically conjugate variables. Each of those quantities are perturbed around a homogeneous and isotropic background.
As usual, the matter sector of the diffeomorphism constraint is uncorrected – as is the Gauss constraint for now.
@@ II.2 Cancellation of the anomalies
Correcting the diffeomorphism constraint leads to a modification of the algebra of constraint and, in particular, to additional anomalies when compared to [18]. Those new anomalies are listed in Appendix A. In order to keep this article as short as possible, we do not make explicit the heavy calculations and we simply focus on the solutions. The approach consists in expressing all the counterterms as functions of the corrections, and then constraining the corrections if any anomalies remain.
The conditions $\mathcal{A}_{2}^{\left\{\mathbb{H},\mathbb{D}\right\}}=0$ and $\mathcal{A}_{7}^{\left\{\mathbb{H},\mathbb{D}\right\}}=0$ (A.60 and A.65) imply $\xi_{1}=\xi_{2}$ as $\alpha_{9}$ and $\alpha_{3}$ cannot be constant so as to ensure a correct classical limit. The same limit considered for $\alpha_{3}$ and $\alpha_{1}$, together with $\mathcal{A}_{5}^{\left\{\mathbb{H},\mathbb{H}\right\}}=0$ (A.80), implies
@@
These counterterms are therefore “not physical” or, if one sees them as mere corrections, not necessary. Then, $\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=0$ (A.54), leads, with Eq. (4), to:
@@
It is important to underline that there are no additional degrees of freedom when compared to [18]. All the counterterms except $\alpha_{9}$ are fixed by the corrections.
@@ III.2 Cancellation of the anomalies
The new list of anomalies is given in Appendix B. This section focuses on the resolution of the associated system.
First, $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.28) implies $\xi_{1}=0$. Combined with $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{D}\}}=0$ and $\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{D}\}}=0$ (B.8 and B.13), the classical limits of $\alpha_{9}$ and $\alpha_{3}$ imply
@@
This causes the anomalies $\mathcal{A}_{8}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{9}^{\{\mathbb{H},\mathbb{D}\}}$ (B.14 and B.15) to vanish immediately.
Considering the classical limits, the cancellation of $\mathcal{A}_{16}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{17}^{\{\mathbb{H},\mathbb{D}\}}$ (B.22 and B.23) gives
@@
leading to $\mathcal{A}_{9}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.32). The cancellations of $\mathcal{A}_{10}^{\{\mathbb{H},\mathbb{D}\}}$, $\mathcal{A}_{11}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{15}^{\{\mathbb{H},\mathbb{D}\}}$ (B.16, B.17 and B.21) imply, respectively,
@@
Adding $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{D}\}}=0$ and $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{D}\}}=0$ (B.18 and B.19) gives $\beta_{10}=\beta_{12}=\beta_{11}$. Re-injecting this result in $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{D}\}}=0$ and subtracting the latter to $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{D}\}}=0$ leads to
@@
Then, in order to be consistent with the classical limit of $\alpha_{3}$, $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.27) implies
@@
Cancelling $\mathcal{A}_{6}^{\{\mathbb{H},\mathbb{D}\}}$ (B.12) then gives
@@
Since $\beta_{4}=0$ (9), $\mathcal{A}_{15}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.38) implies
@@
Re-injecting this result in $\mathcal{A}_{14}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.37), and making use of $\alpha_{4}=\alpha_{5}$, leads to $2\beta_{2}(1+\alpha_{4})=0$. Thus, recovering the correct classical limit for $\alpha_{4}$ imposes $\beta_{2}=0$. Since $2\beta_{1}=\beta_{2}$, this gives
@@
Considering $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.36) leads to
@@
since $\beta_{5}=\beta_{6}$, from Eq. (11). The cancellation of $\mathcal{A}_{17}^{\{\mathbb{H},\mathbb{H}\}}$ (B.40) and $\beta_{1}=0$ imply
@@
Inserting this expression with $\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.30), which implies $\partial_{\mathfrak{c}}\beta_{3}=0$, in $\mathcal{A}_{20}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.43) leads to
@@
This also relies on the fact that $g$ and $\partial_{\mathfrak{c}}g$ cannot vanish, in order to get the correct classical limit ($g\rightarrow\mathfrak{c}$). The term $\beta_{3}$ is then a constant since it depends neither on $\mathfrak{p}$ nor on $\mathfrak{c}$ ($\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{H}\}}=0$). As a consequence, the only way to get the correct classical limit is to have
@@
This cancellation applied to $\mathcal{A}_{14}^{\{\mathbb{H},\mathbb{D}\}}$ (B.20) implies $\beta_{10}=-\beta_{9}$, thus
@@
The cancellation of $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{H}\}}$ (B.35) leads to
@@
The only non-vanishing matter counterterms are thus $\beta_{13}$, which is still undetermined, and $\beta_{10}$, $\beta_{9}$, $\beta_{12}$ and $\beta_{11}$ which are related, by Eq. (23), to the diffeomorphism correction. Those latter counterterms would vanish if the diffeomorphism constraint was not corrected.
At this level, $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{D}\}}=0$ (B.7) implies
@@
Cancelling $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{D}\}}$ (B.9) then implies:
@@
It can be noticed that because of the two previous relations, $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{D}\}}$ (B.10) vanishes without any new condition. Then, injecting the previous expression in $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{D}\}}=0$ (B.11) leads to an expression for $\alpha_{7}$, and thus also for $\alpha_{8}$ :
@@
We next combine the expressions of $\alpha_{1}$, $\alpha_{5}$ and $\alpha_{6}$ in $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.26), then account that cancelling the anomaly $\mathcal{A}_{16}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.39) implies that $\partial_{\mathfrak{c}}\alpha_{3}=0$. This yields to a relation between $\alpha_{3}$ and $\alpha_{9}$,
@@
It is worth noticing that, since $\alpha_{3}$ does not depend on $\mathfrak{c}$, $\alpha_{9}$ does not either. Solving this equation leads to :
@@
As $\alpha_{9}(\mathfrak{p})$ does not appear in any other anomaly and is not constrained in any way, it remains a free function of the theory, which determines, together with the explicit expression of the corrections, $\alpha_{3}$ and thus $\beta_{13}$. All the other counterterms of the Hamiltonian constraint are fully determined by the correction functions and the counterterms from the other constraints. In order to constrain $\xi_{3}$, $\chi_{1}$, $\chi_{2}$, and the corrections, it is important to check the remaining anomalies.
$\mathcal{A}_{1}^{\{\mathbb{D},\mathbb{G}\}}=0$ (B.1) implies
@@
As $\beta_{10}$ and $\beta_{12}$ are related to $(h+\xi_{3})$, if $(l+\chi_{2})$ fulfils Eq. 35, then $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{G}\}}=0$ and $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{G}\}}=0$ (B.5 and B.6) are automatically verified since $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{G}\}}=0$ (B.4) implies $\partial_{\mathfrak{c}}\chi_{1}=0$.
Using this later condition, Eq. (35) and the expressions of $\alpha_{4}$ and $\alpha_{6}$ in $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{G}\}}=0$ (B.2) lead to
@@
with $L$ another integration constant. This allows to re-write Eq. (35) as
@@
Using this in $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{G}\}}=0$ (B.3) directly cancels the anomaly without any new condition on the counter-terms or on the correction function. For $\mathcal{A}_{18}^{\{\mathbb{H},\mathbb{H}\}}$ (B.41), this implies
@@
Finally, $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.24) leads to a restriction on $g$,
@@
which is the same condition as the one obtained with $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{H}\}}=0$ (B.25). Those two anomalies are redundant. This condition is very different from the one obtained in [18, 11]. In particular, Eq. (42) is more general than the condition derived in previous works. Solutions presented in those articles were indeed sufficient to cancel the anomalies. However, if one was looking for the most general solution, Eq. (42) – which is mandatory in our case – would have been found.
@@
This concludes the closure of the algebra for arbitrary correction functions $g$, $\tilde{g}$, $h$ and $l$, as the other anomalies are automatically vanishing with the previous conditions. In summary, the theory requires only one condition on the correction functions given by Eq. (42). Specifically, $l$ and $h$ are unconstrained by the algebra as long as the counterterms are included. Furthermore, there are two new degrees of freedom with respect to the case where only the Hamiltonian constraint is corrected [18].
The first new degree of freedom is the constant $L$ which appears in Eqs. (38) and (39). The latter gives the physical interpretation of $L$: it links the effective Gauss and diffeomorphism corrections at the level of constraints. In particular, it imposes that $(h+\xi_{3})$ must be proportional to $(l+\chi_{2})$ – in other words that the Gauss and diffeomorphism corrections must be the same, up to a proportionality factor.
@@
One should also notice the $\alpha_{9}(\mathfrak{p})$ term which is still a free function of the theory together with the constant $K$, as in [18].
As a summary, the algebra for arbitrary correction functions $g$, $\tilde{g}$, $h$ and $l$ – if the counterterms fulfill the previous relations (see Appendix C) reads:
@@
$\{\mathbb{H}[N_{1}],\mathbb{H}[N_{2}]\}=$ $\frac{1}{2}\partial_{\mathfrak{c}}^{2}g^{2}\left[1+\alpha_{3}\left[K,\alpha_{9}\right](\mathfrak{p})\right]$ (48)

### v2

@@ Abstract
Deforming the algebra of constraint is a well-known approach to effective loop quantum cosmology. More generally, it is a consistent way to modify gravity from the Hamiltonian perspective. In this framework, the Hamiltonian (scalar) constraint is usually the only one to be holonomy corrected. As a heuristic hypothesis, we consider the possibility to also correct the diffeomorphism and Gauss constraints. It is shown that it is impossible to correct the diffeomorphism constraint without correcting the Gauss one, while maintaining a first-class algebra. However, if all constraints are corrected, the algebra can be closed. The resulting differential equations to be fulfilled by the corrections (of the background and of the perturbations) are derived.
@@ I Introduction
In the loop view, quantum geometry corrections are effectively taken into account by replacing the Sen-Ashtekar-Barbero connection by a generalized function that encodes the holonomy along a graph, the so-called holonomy correction. Following recent works – in particular remarks made in [23, 25, 6, 1] – we implement this correction into the classical algebra of general relativity, without specifying its explicit form a priori. Then, we ensure the consistency of the model by requiring the closure of the algebra of constraints – that is, in making sure that each Poisson bracket between constraints is proportional to a constraint, or a combination of constraints: the evolution of a physical solution yields another physical solution. This approach actually extends far beyond its initial purpose and offers a powerful framework for studying modified gravity.
@@
At the end of the day, one should obtain a first-class Poisson algebra and the subtle cancellation game between anomalies and counterterms could lead to restrictions on the allowed forms of the correction functions. This might have observable consequences as the Mukhanov-Sasaki equation for gauge-invariant perturbations is expected to be modified accordingly. Not to mention a possible disappearance of time [9, 4] around the quantum bounce [2].
Usually, the Hamiltonian constraint is the only one to receive corrections. Although there are reasons for this [10], we investigate here, at the heuristic level, the possibility that the diffeomorphism and Gauss constraints may themselves receive corrections. We begin by focusing on the correction of the diffeomorphism constraint alone. We then proceed by considering the full set of corrections affecting all the constraints that define the theory. The differential equations to be fulfilled by the corrections are investigated in detail. We finally conclude with some remarks on possible future studies.
II Correcting only the Hamiltonian and diffeomorphism constraints
@@ II.1 Expression of the constraints
In this work, in addition to the Hamiltonian constraint, the diffeomorphism constraint is also corrected by a function $h(\mathfrak{c},\mathfrak{p})$. Following the usual procedure, three new counterterms $\xi_{i}(\mathfrak{c},\mathfrak{p})$, $i=1,2,3$, are then introduced. The geometrical constraint reads :
@@
with $N^{a}={\bf{N}}^{a}+\delta N^{a}$ the shift function, $E_{i}^{a}=\mathfrak{p}\delta_{i}^{a}+\delta E^{a}_{i}$ the densitized triads, and $K_{a}^{i}=\mathfrak{c}\delta^{i}_{a}+\delta K^{i}_{a}$ their canonically conjugate variables. Each of those quantities are perturbed around a homogeneous and isotropic background.
As usual, the matter sector of the diffeomorphism constraint is uncorrected – as is the Gauss constraint for now. The expressions of each constraint are given in Appendix A.
@@ II.2 Cancellation of the anomalies
Correcting the diffeomorphism constraint leads to a modification of the algebra of constraint and, in particular, to additional anomalies when compared to [16]. Those new anomalies are listed in Appendix B. To keep this article as short as possible, we do not make explicit the heavy calculations and we simply focus on the solutions. The approach consists of expressing all the counterterms as functions of the corrections, and then constraining the corrections if any anomalies remain.
The conditions $\mathcal{A}_{2}^{\left\{\mathbb{H},\mathbb{D}\right\}}=0$ and $\mathcal{A}_{7}^{\left\{\mathbb{H},\mathbb{D}\right\}}=0$ (B.7 and B.12) imply $\xi_{1}=\xi_{2}$ as $\alpha_{9}$ and $\alpha_{3}$ cannot be constant so as to ensure a correct classical limit. The same limit considered for $\alpha_{3}$ and $\alpha_{1}$, together with $\mathcal{A}_{5}^{\left\{\mathbb{H},\mathbb{H}\right\}}=0$ (B.27), implies
@@
These counterterms are therefore “not physical” or, if one sees them as mere corrections, not necessary. Then, $\mathcal{A}_{1}^{\left\{\mathbb{D},\mathbb{G}\right\}}=0$ (B.1), leads, with Eq. (4), to
@@
It is important to emphasize that there are no additional degrees of freedom when compared to [16]. All the counterterms except $\alpha_{9}$ are fixed by the corrections.
@@ III.2 Cancellation of the anomalies
The new list of anomalies is given in Appendix C. This section focuses on the resolution of the associated system.
First, $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.28) implies $\xi_{1}=0$. Combined with $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{D}\}}=0$ and $\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{D}\}}=0$ (C.8 and C.13), the classical limits of $\alpha_{9}$ and $\alpha_{3}$ imply
@@
This causes the anomalies $\mathcal{A}_{8}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{9}^{\{\mathbb{H},\mathbb{D}\}}$ (C.14 and C.15) to vanish immediately.
Considering the classical limits, the cancellation of $\mathcal{A}_{16}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{17}^{\{\mathbb{H},\mathbb{D}\}}$ (C.22 and C.23) gives
@@
leading to $\mathcal{A}_{9}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.32). The cancellations of $\mathcal{A}_{10}^{\{\mathbb{H},\mathbb{D}\}}$, $\mathcal{A}_{11}^{\{\mathbb{H},\mathbb{D}\}}$ and $\mathcal{A}_{15}^{\{\mathbb{H},\mathbb{D}\}}$ (C.16, C.17 and C.21) imply, respectively,
@@
Adding $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{D}\}}=0$ and $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{D}\}}=0$ (C.18 and C.19) gives $\beta_{10}=\beta_{12}=\beta_{11}$. Reinjecting this result in $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{D}\}}=0$ and subtracting the latter from $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{D}\}}=0$ leads to
@@
Then, to be consistent with the classical limit of $\alpha_{3}$, $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.27) implies
@@
Canceling $\mathcal{A}_{6}^{\{\mathbb{H},\mathbb{D}\}}$ (C.12) then gives
@@
Since $\beta_{4}=0$ (9), $\mathcal{A}_{15}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.38) implies
@@
Reinjecting this result in $\mathcal{A}_{14}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.37), and making use of $\alpha_{4}=\alpha_{5}$, leads to $2\beta_{2}(1+\alpha_{4})=0$. Thus, recovering the correct classical limit for $\alpha_{4}$ imposes $\beta_{2}=0$. Since $2\beta_{1}=\beta_{2}$, this gives
@@
Considering $\mathcal{A}_{13}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.36) leads to
@@
since $\beta_{5}=\beta_{6}$, from Eq. (11). The cancellation of $\mathcal{A}_{17}^{\{\mathbb{H},\mathbb{H}\}}$ (C.40) and $\beta_{1}=0$ imply
@@
Inserting this expression with $\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.30), which implies $\partial_{\mathfrak{c}}\beta_{3}=0$, in $\mathcal{A}_{20}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.43) leads to
@@
This also relies on the fact that $g$ and $\partial_{\mathfrak{c}}g$ cannot vanish, to get the correct classical limit ($g\rightarrow\mathfrak{c}$). The term $\beta_{3}$ is then a constant since it depends neither on $\mathfrak{p}$ nor on $\mathfrak{c}$ ($\mathcal{A}_{7}^{\{\mathbb{H},\mathbb{H}\}}=0$). As a consequence, the only way to get the correct classical limit is to have
@@
This cancellation applied to $\mathcal{A}_{14}^{\{\mathbb{H},\mathbb{D}\}}$ (C.20) implies $\beta_{10}=-\beta_{9}$, and, thus,
@@
The cancellation of $\mathcal{A}_{12}^{\{\mathbb{H},\mathbb{H}\}}$ (C.35) leads to
@@
The only nonvanishing matter counterterms are, thus, $\beta_{13}$, which is still undetermined, and $\beta_{10}$, $\beta_{9}$, $\beta_{12}$, and $\beta_{11}$ which are related, by Eq. (23), to the diffeomorphism correction. Those latter counterterms would vanish if the diffeomorphism constraint was not corrected.
At this level, $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{D}\}}=0$ (C.7) implies
@@
Canceling $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{D}\}}$ (C.9) then implies
@@
It can be noticed that because of the two previous relations, $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{D}\}}$ (C.10) vanishes without any new condition. Then, injecting the previous expression in $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{D}\}}=0$ (C.11) leads to an expression for $\alpha_{7}$, and, thus, also for $\alpha_{8}$,
@@
We next combine the expressions of $\alpha_{1}$, $\alpha_{5}$, and $\alpha_{6}$ in $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.26), and then account that canceling the anomaly $\mathcal{A}_{16}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.39) implies that $\partial_{\mathfrak{c}}\alpha_{3}=0$. This yields to a relation between $\alpha_{3}$ and $\alpha_{9}$,
@@
It is worth noticing that, since $\alpha_{3}$ does not depend on $\mathfrak{c}$, $\alpha_{9}$ does not either. Solving this equation leads to
@@
As $\alpha_{9}(\mathfrak{p})$ does not appear in any other anomaly and is not constrained in any way, it remains a free function of the theory, which determines, together with the explicit expression of the corrections, $\alpha_{3}$ and, thus, $\beta_{13}$. All the other counterterms of the Hamiltonian constraint are fully determined by the correction functions and the counterterms from the other constraints. To constrain $\xi_{3}$, $\chi_{1}$, $\chi_{2}$, and the corrections, it is important to check the remaining anomalies.
$\mathcal{A}_{1}^{\{\mathbb{D},\mathbb{G}\}}=0$ (C.1) implies
@@
As $\beta_{10}$ and $\beta_{12}$ are related to $(h+\xi_{3})$, if $(l+\chi_{2})$ fulfills Eq. 35, then $\mathcal{A}_{4}^{\{\mathbb{H},\mathbb{G}\}}=0$ and $\mathcal{A}_{5}^{\{\mathbb{H},\mathbb{G}\}}=0$ (C.5 and C.6) are automatically verified since $\mathcal{A}_{3}^{\{\mathbb{H},\mathbb{G}\}}=0$ (C.4) implies $\partial_{\mathfrak{c}}\chi_{1}=0$.
Using this later condition, Eq. (35) and the expressions of $\alpha_{4}$ and $\alpha_{6}$ in $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{G}\}}=0$ (C.2) lead to
@@
with $L$ another integration constant. This allows one to rewrite Eq. (35) as
@@
Using this in $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{G}\}}=0$ (C.3) directly cancels the anomaly without any new condition on the counterterms or on the correction function. For $\mathcal{A}_{18}^{\{\mathbb{H},\mathbb{H}\}}$ (C.41), this implies
@@
Finally, $\mathcal{A}_{1}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.24) leads to a restriction on $g$,
@@
which is the same condition as the one obtained with $\mathcal{A}_{2}^{\{\mathbb{H},\mathbb{H}\}}=0$ (C.25). Those two anomalies are redundant. This condition is very different from the one obtained in [16, 18]. In particular, Eq. (42) is more general than the condition derived in previous works. Solutions presented in those articles were indeed sufficient to cancel the anomalies. However, if one were looking for the most general solution, Eq. (42) – which is mandatory in our case – would have been found.
@@
This concludes the closure of the algebra for arbitrary correction functions $g$, $\tilde{g}$, $h$, and $l$, as the other anomalies are automatically vanishing with the previous conditions. In summary, the theory requires only one condition on the correction functions given by Eq. (42). Specifically, $l$ and $h$ are unconstrained by the algebra as long as the counterterms are included. Furthermore, there are two new degrees of freedom with respect to the case where only the Hamiltonian constraint is corrected [16].
The first new degree of freedom is the constant $L$ that appears in Eqs. (38) and (39). The latter gives the physical interpretation of $L$: it links the effective Gauss and diffeomorphism corrections at the level of constraints. In particular, it imposes that $(h+\xi_{3})$ must be proportional to $(l+\chi_{2})$ – in other words, that the Gauss and diffeomorphism corrections must be the same, up to a proportionality factor.
@@
One should also notice the $\alpha_{9}(\mathfrak{p})$ term that is still a free function of the theory together with the constant $K$, as in [16].
As a summary, the algebra, at second order in perturbation, for arbitrary correction functions $g$, $\tilde{g}$, $h$ and $l$ – if the counterterms fulfill the previous relations (see Appendix D) reads
@@
$\{\mathbb{H}[N_{1}],\mathbb{H}[N_{2}]\}=$ $\frac{1}{2}\partial_{\mathfrak{c}}^{2}g^{2}\left[1+\alpha_{3}\left[K,\alpha_{9}\right](\mathfrak{p})\right]$